In the period of a year, an investor lost twice as much as he gained. if he originally had $10,000 and now has $8,000. What amount did he gain and what did he lose?
Answer by ankor@dixie-net.com(12701)

In the period of a year, an investor lost twice as much as he gained.
If he originally had $10,000 and now has $8,000.
What amount did he gain and what did he lose?
:
Let x = amt he gained
then
2x = amt he lost
:
$10000 + gained - lost = $8000
;
10000 + x - 2x = 8000
:
10000 - x = 8000
:
-x = 8000 - 10000
:
-x = -2000
:
x = 2000 he gained
and
2x = 4000 he lost